OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y less than or greater to -2x + 10? (0, -4) (1, 13) (6, -1) (3, 8)

OpenStudy (anonymous):

I think your suppose to plug them in :/

OpenStudy (anonymous):

My answer choice is : (3,8)

OpenStudy (nincompoop):

got solution as to how you choose that answer?

OpenStudy (anonymous):

8 <= 2(3) +10 8 <= 6+10 8<=16

OpenStudy (anonymous):

Is that correct ? lol I hope :3

OpenStudy (nincompoop):

you're just basically plugging in... that is not the correct way to solve the problem. assume you didn't have any choices and solve for the answer

OpenStudy (anonymous):

:( whats the way den

OpenStudy (nincompoop):

your solution is incorrect because it is supposed to be y<=-2x+10 if you plug in 3 as your x then your y will be 8 <= -2(3)+10 8 <= -6+10 8 <= 4 therefore (3, 8) is not correct

OpenStudy (anonymous):

(6, -1) ?

OpenStudy (anonymous):

I really think is (0,4) 4<=2(0)+10 4<=10

OpenStudy (nincompoop):

that is correct

OpenStudy (anonymous):

Thank you !!